OpenStudy (anonymous):

Solve the following equation. 8(1 - 5x) = 13 x=

OpenStudy (anonymous):

go to www.wolframalpha.com u will get steps along with your answer :)

OpenStudy (anonymous):

ok so you distribute the 8 first by multiplying it by everything in the brackets. so you'll get 8 - 40x = 13 minus 8 on both sides , -8 + 8 - 40x = 13 - 8 -40x = 5 add 40x on both sides, +40x -40x = 5 + 40x 0 = 5 + 40x minus 5 from both sides, -5 + 0 = 5 + 40x - 5 -5 = 40x and lastly divide both sides by 40, -5/40 = 40x/40 -1/8 = x